# MARS: Optimizing Dual-System Deep Research via Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Large Reasoning Models (LRMs) often exhibit a tendency for overanalysis in simple tasks, where the models excessively utilize System 2-type, deliberate reasoning, leading to inefficient token generation. Furthermore, these models face challenges in adapting their reasoning capabilities to rapidly changing environments due to the static nature of their pre-training data. To address these issues, advancing Large Language Models (LLMs) for complex reasoning tasks requires innovative approaches that bridge intuitive and deliberate cognitive processes, akin to human cognition’s dual-system dynamic. This paper introduces a Multi-Agent System for Deep ReSearch (MARS) enabling seamless integration of System 1’s fast, intuitive thinking with System 2’s deliberate reasoning within LLMs. MARS strategically integrates multiple external tools—such as Google Search, Google Scholar, and Python Interpreter—to access up-to-date information and execute complex computations, while creating a specialized division of labor where System 1 efficiently processes and summarizes high-volume external information, providing distilled insights that expand System 2’s reasoning context without overwhelming its capacity. Furthermore, we propose a multi-agent reinforcement learning framework extending Group Relative Policy Optimization to simultaneously optimize both systems with multi-turn tool interactions, bin-packing optimization, and sample balancing strategies that enhance collaborative efficiency. Extensive experiments demonstrate MARS achieves substantial improvements of 3.86% on the challenging Humanity’s Last Exam (HLE) benchmark and an average gain of 8.9% across 7 knowledge-intensive tasks, validating the effectiveness of our dual-system paradigm for complex reasoning in dynamic information environments.
\$*\$\$*\$footnotetext: Corresponding authors.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across various tasks with their System 1’s fast, intuitive thinking, yet they still struggle with complex reasoning tasks [1, 19, 56, 33, 49]. Recent advances in this direction have led to the emergence of Large Reasoning Models (LRMs), which specifically excel at System 2-type, deliberate reasoning when confronted with challenging problems [20, 15, 50, 2]. However, when faced with simpler questions, LRMs may tend to overanalyze, producing an unnecessary amount of tokens [5]. In contrast to LLMs and LRMs, humans can effortlessly switch between these two modes of thinking.
@@
Furthermore, the pursuit of developing LLMs capable of complex reasoning in rapidly changing environments remains a significant challenge. Since the knowledge these models possess is confined to the cut-off date of their training data, enabling them to acquire new information through web browsing—known as retrieval-augmented generation (RAG)—is emerging as a promising trend for enhancing reasoning capabilities [4, 28, 22, 29]. One of the most prominent examples is the proprietary system [37, 14, 55, 34], the agents designed to synthesize large volumes of online information and accomplish multi-step research tasks effectively.
Inspired by the dual-process theory of human cognition [9, 11], we propose MARS, a multi-agent system for deep research to overcome previous limitations by innovatively integrating System 1 and System 2 thinking mode in a seamless manner, complemented by external tools like Google Search, Google Scholar, and Python Interpreter. This interactive approach allows System 2 to focus on deliberate reasoning and planning, autonomously generating specific queries and computational tasks to engage with external resources. In turn, System 1 processes and summarizes these external tool outputs with its fast, intuitive thinking, distilling high-volume, potentially noisy information into concise insights for System 2’s reasoning processes. The advantages of our proposal are manifest in two critical areas: first, through a specialized division of labor, our method considerably enhances the depth and breadth of available informational context in System 2. Specifically, System 1 efficiently filters and distills large volumes of retrieved information (such as multiple entire research papers or web pages) that would otherwise overwhelm System 2’s reasoning capacity, thereby allowing the model to digest more comprehensive and up-to-date information without sacrificing reasoning quality. Second, this dual-system synergy creates an efficient collaborative framework where each component operates within its optimal domain, resulting in efficient yet robust problem-solving capabilities.
To implement our approach, we develop a data curation pipeline and collect a curated dataset from public datasets that emphasizes diversity across various difficulty levels and academic disciplines. This expansive dataset ensures that the model is well-prepared to tackle a broad spectrum of scenarios, ranging from straightforward tasks to intricate reasoning challenges, creating a solid foundation for subsequent training. To equip LLMs with these dual-system capabilities, we propose a multi-agent reinforcement learning framework that extends the Group Relative Policy Optimization (GRPO) [47] algorithm to optimize System 1 and System 2 simultaneously. In MARS, System 1 and System 2 function as collaborative agents implemented within the same underlying LLM but orchestrated through distinct prompts, enabling seamless interplay between intuitive and deliberate reasoning modes. Additionally, we propose several strategies to optimize our multi-agent training process. First, we incorporate multi-turn and multiple tools use during rollouts, allowing the model to dynamically refine its reasoning through iterative engagement with external tools. Second, we employ bin-packing algorithms [6] to efficiently organize variable-length retrieved content into optimally-sized chunks, significantly enhancing System 1’s parallel processing efficiency. Third, to ensure balanced training dynamics between the two systems, we propose a two-phase approach that pre-computes each system’s advantage signals and then implements a sample balancing strategy between systems, preventing one system from dominating the learning process. By integrating these strategies within our multi-agent RL framework, we enable both systems to adapt and refine their interactions, ultimately enhancing their collaborative efficiency and overall reasoning capabilities. Extensive experiments demonstrate that MARS achieves a substantial improvement of 3.86% on HLE benchmark and an average gain of 8.9% across 7 knowledge-intensive tasks, highlighting the effectiveness of our dual-system paradigm in complex reasoning scenarios.
In summary, our main contributions can be summarized as follows:
  • We propose MARS, a novel dual-system framework that seamlessly integrates System 1’s fast, intuitive thinking mode with System 2’s deliberate reasoning capabilities, creating an efficient collaborative paradigm where each component operates within its optimal domain.
  • We propose a multi-agent reinforcement learning framework to concurrently optimize System 1 and System 2, implementing multi-turn and multiple tools use for System 2, bin-packing optimization for System 1’s efficient content processing, and pre-computation and sampling balance strategies to enhance collaborative efficiency and reasoning capabilities across diverse tasks.
  • We contribute a data curation pipeline and a carefully curated training dataset spanning diverse difficulty levels and academic disciplines, with special emphasis on the challenging Humanity’s Last Exam (HLE) benchmark, for which high-quality training data was previously scarce.
  • We conduct comprehensive experiments on HLE and 7 knowledge-intensive tasks to evaluate the effectiveness of our MARS. Furthermore, we provide in-depth analysis of the multi-agent, multi-turn, multi-tool RL process and investigate how different external tools contribute to reasoning performance, particularly on the challenging HLE benchmark.
@@ 2 Methodology
In this section, we begin by providing a concise overview of the entire pipeline to help readers understand the comprehensive integration of System 1 and System 2 in deep research for tackling complex questions. Then, we delve into the optimization strategies for end-to-end training of the two systems within a multi-agent reinforcement learning framework.
2.1 Dual-System Collaborative Framework for Deep Research
Figure 1: Overview of Dual-System Collaborative Framework in our MARS.
@@ 2.1 Dual-System Collaborative Framework for Deep Research
$\mathcal{P}(\text{answer}|q)=\prod_{i=1}^{N}\left[\underbrace{\pi_{\text{sys}_{2}}(s_{i},t_{i},p_{i}|c_{i})}_{\text{System 2: Reasoning}}\cdot\underbrace{\pi_{\text{sys}_{1}}(\tilde{o}_{t_{i}}|\text{Bin-Packing}(o_{t_{i}}^{(1)},o_{t_{i}}^{(2)},...,o_{t_{i}}^{(n_{t_{i}})}),p_{i})}_{\text{System 1: Information Processing}}\right]$ (4)
@@ 2.2.1 Efficient Content Processing with Bin-Packing
During rollouts, System 2 follows standard token generation for deliberate reasoning and tool-use planning. However, the subsequent tool calls often return multiple outputs of variable length. Processing these large volumes of variable-length text presents a significant challenge for System 1 due to multiple generations. To address this, we employ an efficient bin-packing strategy based on the First Fit Decreasing (FFD) algorithm [6]. This approach optimizes the organization of variable-length content into optimally-sized chunks to reduce the number of generations required by System 1.
Specifically, we begin by counting the number of tokens in each tool output $o_{t_{i}}^{(j)}$. If the token count exceeds System 1’s maximum context length, the output is truncated and placed in an isolated bin. For the remaining outputs, we apply the FFD algorithm: all such outputs are first sorted in decreasing order of their lengths, then assigned to the first bin that can fit them, or to a new bin if no existing one suffices. We choose FFD over Best Fit Decreasing (BFD) due to its superior efficiency in practice. The implementation details can refer to Appendix C.2.
@@ 2.2.2 Advantage pre-computation and Balanced Sampling Mechanism
In our MARS, following HLE’s evaluation prompt [42], we employ LLMs as an evaluator to assess the predicted answer of each trajectory. Note that all System 1 and System 2 samples within the same trajectory share this trajectory-level reward to encourage both systems toward the same goal, rather than pursuing potentially conflicting individual objectives.
For each question $q$, we perform $G$ rollout trajectories and yield exactly $G$ System 2 samples and a variable number of System 1 samples, which depends on the number of tool calls per trajectory and chunks created after Bin-Packing. This unpredictable imbalance can lead to one system dominating the learning process, potentially undermining the collaborative dynamics essential to our multi-agent system. To address this, we pre-compute each system’s advantage and then balance System 1’s samples. Following GRPO [47], rewards are normalized within their corresponding groups to calculate advantage values:
@@ 2.2.3 Multi-Agent Training Objective
For each system, we apply the GRPO objective [47]:
@@
where $\mathcal{D}_{i}$ represents the balanced dataset for System $i$, $x$ and $y$ denote the input-output pairs, $\mathcal{L}_{\text{policy}}$ is the policy loss and $\mathcal{L}_{\text{KL}}$ is a KL regularization term. The detailed mathematical formulations are provided in Appendix C.4. This joint optimization approach allows both systems to improve simultaneously while maintaining their specialized roles in the collaborative framework.
@@ 3.1 Evaluation Datasets and Metrics
We evaluate our MARS on several challenging benchmarks that require sophisticated multi-step reasoning and external knowledge: (1) Humanity’s Last Exam (HLE) [42] is an extremely challenging dataset containing advanced problems. We utilize its text-only subset with 2,154 questions. (2) Single-Hop Question Answering, including NQ [24], TriviaQA [23], and PopQA [32]. (3) Multi-Hop Question Answering, including HotpotQA [58], 2WikiMultiHopQA [17], Musique [51], and Bamboogle [43]. For evaluation, we follow different protocols for different benchmarks. For HLE, we adopt its official evaluation prompt [42] with GPT-4o [19] as the judge. For other QA datasets, following [4], we employ Qwen2.5-72B-Instruct [56] as the evaluation model.
3.2 Baselines and Implementation Details
To evaluate the effectiveness of MARS, we compare our method with the following baselines. (1) Direct Reasoning: Models that directly answer questions without external knowledge, including open-source models (Qwen2.5 series [56] and QwQ-32B [50]) and powerful proprietary models (DeepSeek-R1-671B [15], GPT-4o [19], o1 [36], Claude 3.7 Sonnet [2], Gemini 2.5 Pro [13], o3(high) [38]), and o4-mini(high) [39]. (2) Advanced RAG Reasoning Methods: We consider the standard RAG that retrieve top-10 documents based on the question, and several iterative RAG methods, including Self-RAG [3], InstructRAG [53], Auto-RAG [60] and C-3PO [4]. (3) R1-like Reasoning with Search: Methods that integrate external knowledge into R1-like reasoning, including Search-o1 [28], Search-R1 [22], WebThinker [29], and OpenAI Deep Research [37].
We initialize our policy model with Qwen2.5-7B-Instruct [56] and Qwen3-8B [57] and design dedicated prompts for System 1 and System 2 (Appendix D) to facilitate multi-agent reinforcement learning. Additional implementation details are provided in Appendix C.
3.3 Main Results on Humanity’s Last Exam
Table 1: Main Results on HLE (evaluated with official evaluation prompt [42] by GPT-4o). Results for proprietary models are from the official leaderboard for reference. For open-source models, the best results are in bold and the second are underlined.
@@ 3.3 Main Results on Humanity’s Last Exam
Table 1 compares the performance of our MARS with various baselines on the Humanity’s Last Exam (HLE) benchmark. This benchmark is particularly challenging as it contains advanced problems across multiple disciplines that require sophisticated reasoning and up-to-date knowledge. We have observed several key findings:
First, MARS achieves an average accuracy of 7.38% across all categories, representing a substantial improvement of 3.86% over the base model (Qwen2.5-7B-Instruct). Notably, MARS outperforms all other open-source models and reasoning methods, including those based on much larger models such as WebThinker(QwQ-32B) and C-3PO(Qwen2.5-72B-Instruct). This demonstrates that our dual-system paradigm effectively enhances reasoning capabilities with significantly fewer parameters, achieving superior performance. Second, when examining subject-specific performance, MARS demonstrates particularly strong results in knowledge-intensive domains. This advantage stems from our dual-system design, where System 1 efficiently processes and distills large volumes of retrieved information from web pages and research papers without consuming System 2’s reasoning tokens. This architecture allows System 2 to focus exclusively on complex reasoning while having access to comprehensive, filtered knowledge, which is especially beneficial for technical domains requiring both extensive background information and multi-step reasoning. Finally, while proprietary models like OpenAI Deep Research still maintain a lead in overall performance, MARS significantly narrows the gap between open-source and commercial solutions. The performance gaps between our MARS (7.38%) and proprietary models like Claude 3.7 Sonnet (7.89%) or o1 (7.75%) is notably small, especially considering that MARS utilizes only a 7B parameter model. These results demonstrate that our dual-system paradigm creates an effective synergy: System 1 efficiently filters and distills large volumes of external information, while System 2 maintains focused, deliberate reasoning without token consumption trade-offs.
3.4 Main Results on Knowledge-intensive Reasoning
@@
We present the results on 7 knowledge-intensive question answering tasks in Table 2. The evaluation covers both single-hop (NQ, TriviaQA, PopQA) and multi-hop (HotpotQA, 2Wiki, Musique, Bamboogle) datasets, revealing several key findings:
First, our MARS consistently outperforms all baseline methods across all benchmarks, achieving an impressive improvement of 8.95% over the previous state-of-the-art method C-3PO. This consistent performance gain across diverse datasets demonstrates the robustness and generalizability of our dual-system paradigm for knowledge-intensive reasoning tasks. Second, the performance gains are particularly pronounced on multi-hop reasoning tasks, where MARS achieves an average improvement of 12.2% over C-3PO across the four multi-hop benchmarks. This significant enhancement demonstrates that our approach excels at complex reasoning chains requiring multiple steps of information retrieval and integration. The multi-agent framework allows System 2 to decompose complex questions into manageable sub-queries, while System 1 efficiently processes retrieved information, creating an effective synergy for multi-hop reasoning. Third, even on relatively simpler single-hop QA tasks, MARS achieves an average gain of 7.5% over the previous best methods. This indicates that our dual-system approach enhances performance across the entire spectrum of reasoning complexity, from straightforward factual queries to intricate multi-step problems. By optimizing the collaboration between System 1’s efficient information processing and System 2’s deliberate reasoning capabilities, our MARS effectively leverages external knowledge sources while maintaining computational efficiency. These results further confirm that our approach mitigates the limitations of previous methods by creating a more balanced and effective division of labor between intuitive and deliberate reasoning processes.
3.5 Analysis of Multi-agent RL Process

### v2

# MARS: Co-evolving Dual-System Deep Research via Multi-Agent Reinforcement Learning

@@ Abstract
Large Reasoning Models (LRMs) face two fundamental limitations: excessive token consumption when overanalyzing simple information processing tasks, and inability to access up-to-date knowledge beyond their training data. We introduce MARS (Multi-Agent System for Deep ReSearch), a novel co-evolution framework that jointly optimizes dual cognitive systems through multi-agent reinforcement learning. Unlike prior approaches that employ fixed or independently-trained summarizers, MARS enables System 1 (fast, intuitive processing) and System 2 (deliberate reasoning) to co-adapt through shared trajectory rewards, developing complementary strategies where System 1 learns to distill information specifically useful for System 2’s reasoning. We extend Group Relative Policy Optimization (GRPO) for multi-agent settings with three key innovations: (1) decoupled gradient computation ensuring proper credit assignment despite shared rewards, (2) bin-packing optimization for efficient parallel information processing, and (3) advantage-weighted balanced sampling preventing training imbalance. Extensive experiments demonstrate that MARS (8B), trained under a challenging Zero RL setting without any supervised fine-tuning, achieves 8.17% on HLE—outperforming WebThinker (32B with SFT, 6.87%) and narrowing the gap with proprietary models like Claude 3.7 Sonnet (7.89%)—while achieving an average gain of 8.9% across 7 knowledge-intensive tasks.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across various tasks with their System 1’s fast, intuitive thinking, yet they still struggle with complex reasoning tasks (Anthropic, 2024; Hurst et al., 2024; Yang et al., 2024; Mesnard et al., 2024; Team, 2025). Recent advances in this direction have led to the emergence of Large Reasoning Models (LRMs), which specifically excel at System 2-type, deliberate reasoning when confronted with challenging problems (Jaech et al., 2024; Guo et al., 2025; Team, 2024; Anthropic, 2025). However, when faced with simpler information processing tasks such as analyzing web pages or summarizing papers, LRMs tend to overanalyze, consuming excessive reasoning tokens that limit their capacity to process large volumes of external information (Chen et al., 2024c). In contrast, humans effortlessly switch between fast, intuitive processing and slow, deliberate reasoning—a flexibility that current LRMs lack.
@@
Furthermore, developing LLMs capable of complex reasoning in rapidly changing environments remains a significant challenge. Since model knowledge is confined to the cut-off date of training data, enabling access to external information through retrieval-augmented generation (RAG) is emerging as a promising direction (Chen et al., 2025a; Li et al., 2025a; Jin et al., 2025; Li et al., 2025b; Qiao et al., 2025). State-of-the-art proprietary systems (OpenAI, 2025a; Google, 2025a; xAI, 2025; MoonshotAI, 2025; Team et al., 2025) demonstrate the potential of agents that synthesize large volumes of online information for multi-step research tasks.
Recent open-source efforts such as WebThinker (Li et al., 2025b) and Search-R1 (Jin et al., 2025) have attempted to combine reasoning models with external tools. However, these approaches typically employ fixed or independently-trained information processors that are not optimized for the downstream reasoning task. For example, WebThinker uses a fixed extraction prompt to process retrieved documents, while Search-R1 directly feeds raw search results to the reasoning model. This architectural limitation creates a fundamental mismatch: the information processor cannot learn what specific details the reasoning model needs, leading to either information overload or critical detail loss.
Inspired by the dual-process theory of human cognition (Evans and Stanovich, 2013; Frankish, 2010), we propose MARS (Multi-Agent System for Deep ReSearch), a novel framework that enables co-evolution between dual cognitive systems through multi-agent reinforcement learning. Unlike prior work, MARS jointly optimizes System 1 (fast, intuitive information processing) and System 2 (deliberate reasoning) through shared trajectory rewards, allowing them to develop complementary strategies. This co-evolutionary design is motivated by a key insight: the optimal information distillation strategy depends on the reasoning task, and both systems must co-adapt to achieve effective collaboration. Through joint optimization, System 1 learns to extract information that is specifically useful for System 2’s reasoning—not generic summaries, but task-relevant insights that emerge from their collaborative training.
The advantages of our co-evolution framework are twofold. First, through learned specialization, System 1 efficiently filters and distills large volumes of retrieved information (multiple web pages or research papers per query) that would otherwise overwhelm System 2’s context capacity, enabling the model to digest comprehensive, up-to-date information without sacrificing reasoning quality. Second, this dual-system synergy creates an efficient collaborative framework where each component operates within its optimal domain—System 1 handles high-throughput information processing while System 2 focuses on complex multi-step reasoning—resulting in both efficiency and robustness.
To implement our approach, we extend Group Relative Policy Optimization (GRPO) (Shao et al., 2024) for multi-agent settings with three key innovations. First, we design a decoupled gradient computation mechanism where System 1 and System 2 receive gradients from non-overlapping token sets despite sharing trajectory-level rewards, ensuring proper credit assignment. Second, we employ bin-packing optimization to efficiently organize variable-length retrieved content into optimally-sized chunks, enabling parallel processing by System 1. Third, we propose advantage-weighted balanced sampling that pre-computes advantages for all samples before balancing, preventing either system from dominating the learning process while preserving statistical integrity. We also develop a data curation pipeline to collect diverse, graduate-level training examples from public sources, addressing the scarcity of high-quality data for challenging benchmarks like HLE.
Importantly, MARS operates under a challenging “Zero RL” setting: we train directly from the base model without any Supervised Fine-Tuning (SFT) or policy distillation from stronger models. This contrasts with prior work such as WebThinker (which uses SFT on curated trajectories) and enables direct evaluation of our multi-agent RL framework’s effectiveness. Experiments show that MARS reaches 8.17% accuracy on Humanity’s Last Exam (HLE) with an 8B model, surpassing WebThinker (32B with SFT, 6.87%) and approaching Claude 3.7 Sonnet (7.89%). Across 7 knowledge-intensive QA tasks, it improves performance by 8.9% on average. Ablations further show that removing System 1 reduces HLE accuracy by 1.91%, indicating that System 1 learns task-relevant distillation rather than acting as a passive summarizer.
In summary, our main contributions are:
  • We introduce MARS, a co-evolution dual-system framework that jointly trains System 1 (information processing) and System 2 (reasoning) with shared trajectory rewards, enabling task-specific information distillation.
  • We adapt GRPO to multi-agent RL with three techniques: decoupled gradients for credit assignment, bin-packing for parallel information processing, and advantage-weighted balanced sampling for stable joint optimization.
  • Under a Zero RL setting (no SFT), MARS (8B) achieves 8.17% on HLE and improves 7 knowledge-intensive QA tasks; ablations confirm the benefit of co-evolution and each component.
2 Related Work
Retrieval-Augmented Generation (RAG).
RAG has emerged as a crucial approach to overcome knowledge limitations of large language models by integrating external information sources (Lewis et al., 2020; Nakano et al., 2021; Schick et al., 2023; Yu et al., 2024c; Asai et al., 2024; Jiang et al., 2024; Wei et al., 2025; Yu et al., 2024b; Chen et al., 2025a; Li et al., 2025b). However, existing RAG systems face a fundamental dilemma: they either suffer from information overload when processing multiple lengthy documents (such as entire web pages or research papers), or lose critical details when aggressively condensing information (Fan et al., 2024; Gao et al., 2023; Tan et al., 2024). Our approach addresses this through a dual-system framework where System 2 handles complex reasoning while System 1 efficiently processes retrieved information, enabling MARS to manage larger volumes of external information while maintaining reasoning quality.
@@ 3 Methodology
This section first overviews the pipeline, highlighting how System 1 and System 2 collaborate for deep research. We then present our optimization strategies for end-to-end training under a multi-agent reinforcement learning framework.
3.1 Dual-System Collaborative Framework
Figure 1: Overview of Dual-System Collaborative Framework.
@@ 3.1 Dual-System Collaborative Framework
$\begin{split}\mathcal{P}(&\text{answer}|q)=\prod_{i=1}^{N}\bigg[\underbrace{\pi_{\text{sys}_{2}}(s_{i},t_{i},p_{i}|c_{i})}_{\text{System 2: Reasoning}}\\ &\cdot\underbrace{\pi_{\text{sys}_{1}}(\tilde{o}_{t_{i}}|\text{Bin-Packing}(o_{t_{i}}^{(1)},\ldots,o_{t_{i}}^{(n_{t_{i}})}),p_{i})}_{\text{System 1: Information Processing}}\bigg]\end{split}$ (4)
@@ 3.2.1 Efficient Content Processing with Bin-Packing
During rollouts, System 2 follows standard token generation for deliberate reasoning and tool-use planning. However, the subsequent tool calls often return multiple outputs of variable length. Processing these large volumes of variable-length text presents a significant challenge for System 1 due to multiple generations. To address this, we employ an efficient bin-packing strategy based on the First Fit Decreasing (FFD) algorithm (Coffman Jr et al., 1984). This approach optimizes the organization of variable-length content into optimally-sized chunks to reduce the number of generations required by System 1.
Specifically, we begin by counting the number of tokens in each tool output $o_{t_{i}}^{(j)}$. If the token count exceeds System 1’s maximum context length, the output is truncated and placed in an isolated bin. For the remaining outputs, we apply the FFD algorithm: all such outputs are first sorted in decreasing order of their lengths, then assigned to the first bin that can fit them, or to a new bin if no existing one suffices. We choose FFD over Best Fit Decreasing (BFD) due to its superior efficiency in practice.
@@ 3.2.2 Advantage pre-computation and Balanced Sampling Mechanism
In our MARS, following HLE’s evaluation prompt (Phan et al., 2025), we employ LLMs as an evaluator to assess the predicted answer of each trajectory. Note that all System 1 and System 2 samples within the same trajectory share this trajectory-level reward to encourage both systems toward the same goal, rather than pursuing potentially conflicting individual objectives.
For each question $q$, we perform $G$ rollout trajectories and yield exactly $G$ System 2 samples and a variable number of System 1 samples, which depends on the number of tool calls per trajectory and chunks created after Bin-Packing. This unpredictable imbalance can lead to one system dominating the learning process, potentially undermining the collaborative dynamics essential to our multi-agent system. To address this, we pre-compute each system’s advantage and then balance System 1’s samples. Following GRPO (Shao et al., 2024), rewards are normalized within their corresponding groups to calculate advantage values:
@@ 3.2.3 Multi-Agent Training Objective
For each system, we apply the GRPO objective (Shao et al., 2024):
@@
where $\mathcal{D}_{i}$ represents the balanced dataset for System $i$, $x$ and $y$ denote the input-output pairs, $\mathcal{L}_{\text{policy}}$ is the policy loss and $\mathcal{L}_{\text{KL}}$ is a KL regularization term.
Credit Assignment via Decoupled Gradients. A natural question arises: how can System 1 learn useful summarization behavior when it shares the same trajectory-level reward as System 2? We address this through decoupled gradient computation, which ensures proper credit assignment despite shared rewards:
  • Non-overlapping token sets: System 2’s loss is computed only over reasoning and planning tokens $\{s_{i},t_{i},p_{i}\}$, with System 1’s output tokens $\tilde{o}_{t_{i}}$ masked during loss computation. System 1’s loss is computed only over its distillation outputs, conditioned on the bin-packed input and purpose.
  • Group-specific advantage normalization: Each system’s advantage (Eq. 5-6) is normalized within its own sample group. This means System 1 learns “what made my summary better than other summaries in this group?” rather than simply “was the overall trajectory successful?”
Together, the shared reward aligns both systems toward correct problem solving, while decoupled gradients provide role-specific learning signals: System 2 improves planning and reasoning, and System 1 improves the utility of its distillation. This enables co-evolution: System 2 learns better purposes $p_{i}$, and System 1 learns more task-relevant extractions.
@@ 4.1 Evaluation Datasets and Metrics
We evaluate our MARS on several challenging benchmarks that require sophisticated multi-step reasoning and external knowledge: (1) Humanity’s Last Exam (HLE) (Phan et al., 2025) is an extremely challenging dataset containing advanced problems. We utilize its text-only subset with 2,154 questions. (2) Single-Hop Question Answering, including NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and PopQA (Mallen et al., 2023). (3) Multi-Hop Question Answering, including HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), Musique (Trivedi et al., 2022), and Bamboogle (Press et al., 2023). For evaluation, we follow different protocols for different benchmarks. For HLE, we adopt its official evaluation prompt (Phan et al., 2025) with GPT-4o (Hurst et al., 2024) as the judge. For other QA datasets, following (Chen et al., 2025a), we employ Qwen2.5-72B-Instruct (Yang et al., 2024) as the evaluation model.
Table 1: Comparison of training configurations across methods. MARS achieves superior results despite using smaller models and no SFT warmup.
Method Model Size SFT RL Co-evolution
WebThinker 32B
C-3PO 72B
Search-R1 7B
Search-o1 7B
MARS (Ours) 7B/8B
Table 2: Main Results on HLE (evaluated with official evaluation prompt (Phan et al., 2025) by GPT-4o). Results for proprietary models are from the official leaderboard for reference. For open-source models, the best results are in bold and the second-best are underlined. MARS operates under Zero RL (no SFT) with smaller models yet outperforms larger SFT-trained baselines.